MICROWAVE OBSERVATION OF THE OH-H_{2}O RADICAL COMPLEX

\underline{CAROLYN S. BRAUER}, GALEN SEDO, ERIK M. GRUMSTRUP and KENNETH R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455; MARK D. MARSHALL and HELEN O. LEUNG, Department of Chemistry, Amherst College, Amherst, MA 01002-5000.

The radical complex OH-H_{2}O has been observed by rotational spectroscopy. Spectra for ^{1}^{6}OH-^{1}^{6}OH_{2}O and ^{1}^{8}OH-^{1}^{8}OH_{2} have been analyzed using a two-state model which accounts for nuclear motion on both the ^{2}A′ and ^{2}A′′ potential surfaces. Partial quenching of the OH orbital angular momentum dramatically affects the rotational spectra, and the^{2}A′-^{2}A′′ energy separation, ρ, is determined to be -146.50744(42) cm^{−1}. The ground state of the complex has approximately 86% ^{2}A′ character and the vibrationally averaged OH-OH_{2} hydrogen bond distance is 1.952 Å. The magnetic hyperfine constants for the OH proton in the complex are significantly altered from monomer values.